The response to a question has three alternatives: A, B, and C. A sample of 120 responses provides 58 A, 25 B, and 37 C responses. Show the frequency and relative frequency distributions (use nearest whole number for the frequency column and 2 decimal for the relative frequency column). Class Frequency Relative Frequency A B C (Total)
Answer #1

Before answering, let me define frequency and relative frequency:

1: Frequency : Total number of time an event has occurred.

2: Relative Frequency: A relative frequency is the fraction of times an event occurs.

Relative frequency= No of times an event occurred/Sample Size

Relative frequency for response A=58/120=0.48

Relative frequency for response B=25/120=0.21

Relative frequency for response C=37/120=0.31

Response Frequency Relative Frequency
A 58 0.48
B 25 0.21
C 37 0.31
Total 120 1

Hope this answers your question.

Regards.